# Performance Gap in Entity Knowledge Extraction Across Modalities in Vision Language Models

### v1

@@ Abstract
Vision-language models (VLMs) excel at extracting and reasoning about information from images. Yet, their capacity to leverage internal knowledge about specific entities remains underexplored. This work investigates the disparity in model performance when answering factual questions about an entity described in text versus depicted in an image. Our results reveal a significant accuracy drop—averaging 19%—when the entity is presented visually instead of textually. We hypothesize that this decline arises from limitations in how information flows from image tokens to query tokens. We use mechanistic interpretability tools to reveal that, although image tokens are preprocessed by the vision encoder, meaningful information flow from these tokens occurs only in the much deeper layers. Furthermore, critical image processing happens in the language model’s middle layers, allowing few layers for consecutive reasoning, highlighting a potential inefficiency in how the model utilizes its layers for reasoning. These insights shed light on the internal mechanics of VLMs and offer pathways for enhancing their reasoning capabilities.
@@
Figure 1: Illustration of our motivating observation. The model successfully extracts its knowledge about the entity when the entity is represented in text. It further successfully identifies the entity in the image. But it fail to combine the two.
@@ 1 Introduction
Over the past several years, vision-language models (VLMs) have made remarkable progress in linking visual inputs with natural language descriptions, enabling them to perform tasks such as image recognition Radford et al. (2021); Mirza et al. (2024), video understanding Lin et al. (2024), and Visual Question Answering (VQA) Liu et al. (2023b, a). Much of the existing research evaluating these models focuses on their capacity to understand and describe information presented in the input image: counting objects, assessing complex visual arrangements, and understanding interactions between humans and objects Dang et al. (2024); Basu et al. (2024). While this has demonstrated that VLMs can integrate visual and textual information effectively, it does not fully address a critical question:
@@
how well can VLMs exploit their internal knowledge about entities presented in an image rather than simply described them in text?
@@
This question opens a new angle on VLM evaluation. Instead of probing their grasp of broad scenes, we consider a different scenario: retrieving factual knowledge about visually presented entities. For example, if the subject of an image is a notable figure, say Robbin Williams, we can ask the model "Who is Robbin William’s spouse?", which we refer to as a textual cue. Or we can provide the model with an image of Robbin Williams and ask “Who is the spouse of the subject of this image?”, which we refer to as a visual cue. The second case can be viewed as a two-hop reasoning task. First, the system must recognize the entity visually—identification. Next, it must connect this recognized entity to its stored factual knowledge—extraction. Figure 1 provides an example where the model contains the knowledge to answer the question, but fails to extract it when provided with a visual representation. A detailed explanation about this evaluation appears in section 2.
@@
Our empirical findings reveal a large accuracy drop when relying on visual cues rather than textual cues. This likely stems from the model investing too many layers in identifying the subject visually, leaving fewer layers for the second hop of factual reasoning (Biran et al., 2024). To investigate these dynamics more concretely, we utilize techniques of mechanistic interpretability such as attention knockout (Geva et al., 2023), activation patching (Vig et al., 2020) and vocabulary projection (nostalgebraist, 2020; Geva et al., 2021). They allow us to intervene in the model’s processing mid-inference, altering how and when image information influences the final answer. By tracking the flow of information through the network, we show that entity recognition and knowledge extraction are deeply entangled processes, with identification dominating the representational space longer than what one would expect. Consequently, the model’s internal knowledge is leveraged less efficiently.
@@
The contribution of our work is providing a better understanding of how VLMs process entities and integrate their internal factual knowledge with image information, highlighting a gap between the processing of visual and textual information. Instead of treating entity-level knowledge extraction as a monolithic process, we reveal that it involves at least two interdependent steps. We show that the initial reliance on visual cues can restrict later factual reasoning compared to textual cues. This understanding highlights the need for more multi-modality specific evaluation protocols. Our experiments, which measure the performance gap and apply attention-based hidden state interventions, provide a foundation for future efforts to enhance the models’ ability to draw on their underlying knowledge when presented with visual inputs.
We provide next, a motivational experiment pointing to a gap between textual and visual representations. Then we present our empirical approach to further analyze it. Related work is deferred to Appendix 4. We release our code and data at https://github.com/ido-co/vlm-modality-gap.
2 Motivation
We start by quantifying the model’s performance gap in entity knowledge extraction when presented with visual versus textual representations of entities. This experiment motivates the subsequent experiments which aim to explore and explain this gap. We use Llava-1.5-7B (Liu et al., 2023b) as the VLM in all experiments.
Experiment.
We checked entity recognition by prompting the model to identify the subject in an image and comparing its response to a list of aliases for that entity, aggregated from Wikidata (Vrandečić and Krötzsch, 2014). Using the IMDB-WIKI dataset (Rothe et al., 2015), we sampled 10,000 entity-image pairs and retained the 300 where the model successfully recognized the entity by prompting the model with an image and the prompt “Identify by name the subject of this image”. This step is necessary to isolate the knowledge extraction capabilities of the model from its ability to recognize entities in images.
@@ 2 Motivation
In the textual input case, the model is presented with the question as is. For the visual input, the model receives an image of the entity along with a reformulated query, such as “Where was the subject in the image born?”. This ensures that the model needs to internally identify the subject depicted in the image before reasoning and generating an answer.
@@ Results.
We measured accuracy per entity as the proportion of correctly answered questions for that entity and calculated the overall mean accuracy across all entities. The mean accuracy for textual inputs is 52%, compared to 38% for visual inputs. The average per-entity accuracy drop is 19.5%.
@@
Interestingly, a common mistake made by the model when asked to name the subject’s spouse/parent/sibling/child, is to name the subject itself. This observation leads to our the hypothesis that by the time information about the entity propagates from the image tokens to later token positions, there is not enough computational capacity left to use this information, and the model outputs the entity’s name.
@@
Some entities did exhibit an increase in accuracy with visual inputs. This specifically occurred when the textual input alone was insufficient for reasoning, but the visual input provided contextually inferable details. For instance, a subject wearing a French national team uniform implicitly informed the model that the language spoken by the subject is French.
3 Experiments
We aim to understand at what point in the inference pass the entity in an image is identified, and this identity is propagated to the query and generated token positions. To this end, we will apply interventions in the model’s inference. We continue to focus on entity-image pairs for which, without intervention, the model correctly answers the prompt "Identify by name the subject of this image."
Notation.
Following Liu et al. (2023b), denote by $\mathbf{X}_{\mathbf{v}}$ an input image, and by $\mathbf{H}_{\mathbf{v}}=\mathbf{W}\cdot g(\mathbf{X}_{\mathbf{v}})$ its visual features extracted by the CLIP vision encoder and projected to a sequence of language embedding tokens.
For a text query, $\mathbf{X}_{\mathbf{q}}$, let $\mathbf{H}_{\mathbf{q}}$ denote the sequence of language embedding tokens produced by applying the language tokenizer to $\mathbf{X}_{\mathbf{q}}$.
The input to the language model, with hidden dimension $d$, is formed by concatenating the token sequence $\mathbf{H}_{\mathbf{v}}$ with $\mathbf{H}_{\mathbf{q}}$. Let $\mathbf{h}_{i,v}^{\ell}$ and $\mathbf{h}_{i,q}^{\ell}$ be the hidden representation for the $i^{\text{th}}$ vision and query language tokens. As the language model processes the entire input at each layer $l$, it builds a representation $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell},\mathbf{h}_{1,q}^{\ell}\dots\mathbf{h}_{m,q}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}\right)$, where $\mathbf{h}_{k,g}^{\ell}$ are the generated tokens hidden representations, $\left|\mathbf{H}_{\mathbf{v}}\right|=n$, $\left|\mathbf{H}_{\mathbf{q}}\right|=m$ and $\ell=0$ refers to the initial token embeddings inputted to the LLM.
In our experiments we focus on how information flows from the vision hidden states $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)$ to the query and generated tokens. We do this by applying techniques of mechanistic interpretability.
Experiments Overview.
In the experiments, we begin in Section 3.1 by locating an upper bound on where the model extracts the identity from the image, using activation patching (Vig et al., 2020) between different entities. Having established an upper bound, we ask if the layers up to this bound are efficiently utilized, or if the model could have extracted the identity sooner. We study this in Section 3.2 using forward activation patching. We defer to the appendix an additional experiment which uses attention knockouts (Geva et al., 2023) and complements the results of Section 3.1.
3.1 Late information flow
@@
We begin by identifying the point of critical information flow from the the image hidden representation to later positions. This will provide an upper bound on the layer depth at which entity identification happens and point us to where this information is likely to propagate onward. We can subsequently infer that by this point the model mostly finalizes its image-based representation enrichment and focuses on leveraging this information for prediction. We hypothesize that this happens in deep layers, leaving few layers for factual knowledge extraction, which may begin to explain the performance gap presented in Section 2.
@@
The experiment involves two distinct forward passes through the model, using different entities: an original entity and a injected entity. Each forward pass is initiated with a prompt asking the model to identify the subject of an image. We extract the image-token representations $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$ from the injected entity forward pass at a specific layer. These representations are then patched at the same token positions and same layer of the forward pass of the original entity. After patching, we allow the model to complete its forward pass and predict which entity is identified. The layer at which the predicted entity switches from being the injected entity to the original entity identifies the critical point where the model has completed sufficient processing of visual cues such that patching does not affect its prediction.

### v2

@@ Abstract
Vision-language models (VLMs) excel at extracting and reasoning about information from images. Yet, their capacity to leverage internal knowledge about specific entities remains underexplored. This work investigates the disparity in model performance when answering factual questions about an entity described in text versus depicted in an image. Our results reveal a significant accuracy drop — reaching 18% for some models — when the entity is presented visually instead of textually. To study this gap we present POPVQA, a dataset which allows separating entity recognition and question answering, and use it to benchmark several models. We hypothesize that this decline arises from limitations in how information flows from image tokens to query tokens. Thus, we use mechanistic interpretability tools to reveal that, although image tokens are preprocessed by the vision encoder, meaningful information flow from these tokens occurs only in the much deeper layers. Furthermore, critical image processing happens in the language model’s middle layers, allowing few layers for consecutive reasoning, highlighting a potential inefficiency in how the model utilizes its layers for reasoning. These insights shed light on the internal mechanics of VLMs and offer pathways for enhancing their reasoning capabilities. POPVQA can be found at this link.
@@
Figure 1: Illustration of our motivating observation. The model successfully extracts its knowledge about the entity when the entity is represented in text. It further successfully identifies the entity in the image, but it fails to combine the two.
@@ 1 Introduction
Over the past several years, vision-language models (VLMs) have made remarkable progress in linking visual inputs with natural language descriptions, enabling them to perform tasks such as image recognition Radford et al. (2021a); Mirza et al. (2024), video understanding Lin et al. (2024), and Visual Question Answering (VQA) Liu et al. (2023b, a). Much of the existing research evaluating these models focuses on their capacity to understand and describe information presented in the input image: counting objects, assessing complex visual arrangements, and understanding interactions between humans and objects Dang et al. (2024); Basu et al. (2024). While this has demonstrated that VLMs can integrate visual and textual information effectively, it does not fully address a critical question:
@@
how well can VLMs exploit their internal knowledge about entities presented in an image rather than ones simply described in text?
@@
This question opens a new angle on VLM evaluation. Instead of probing their grasp of broad scenes, we consider a different scenario: retrieving factual knowledge about visually presented entities. For example, if the subject of a question is a notable figure, say Robin Williams, we can ask the model "Who is Robin William’s spouse?", which we refer to as a textual representation. Or we can provide the model with an image of Robin Williams and ask “Who is the spouse of the subject in this image?”, which we refer to as a visual representation. The second case may be viewed as a two-hop reasoning task. First, the model must recognize the entity visually—identification. Next, it must connect this recognized entity to its stored factual knowledge—extraction. Figure 1 provides an example where the model contains the knowledge to answer the question, but fails to extract it when provided with a visual representation. A detailed explanation of this evaluation appears in Section 4.
@@
Our empirical findings reveal a large accuracy drop when relying on visual representations rather than textual ones. Further probing suggests that this stems from the model investing too many layers in identifying the subject visually, leaving fewer layers for the second hop of factual reasoning (Biran et al., 2024). To investigate these dynamics more concretely, we utilize techniques of mechanistic interpretability such as activation patching (Vig et al., 2020) and attention knockout (Geva et al., 2023). They allow us to intervene in the model’s processing mid-inference, altering how and when image information influences the final answer. By tracking the flow of information through the network, we show that entity recognition and knowledge extraction are deeply entangled processes, with identification dominating the representational space longer than what one would expect. Consequently, the model’s internal knowledge is leveraged less efficiently.
@@
The contribution of our work is twofold. First, we release POPVQA, a dataset that facilitates validating entity identification prior to question answering. Second, we provide a better understanding of how VLMs process entities and integrate their internal factual knowledge with image information, highlighting a gap between the processing of visual and textual information. Instead of treating entity-level knowledge extraction as a monolithic process, we reveal that it involves at least two interdependent steps. We show that the initial reliance on visual representations can restrict later factual reasoning compared to textual representations. This understanding highlights the need for more multi-modality specific evaluation protocols.
Our experiments, which measure the performance gap and apply activation-based hidden state interventions, provide a foundation for future efforts to enhance the models’ ability to draw on their underlying knowledge when presented with visual inputs. Code to reproduce our experiments can be found at this link.
2 Problem Setup
Goal.
We ask how well VLMs can leverage their internal factual knowledge when presented with visual representations of entities, compared to textual ones. We suggest that VLMs perform two key steps: identifying the entity in the image, followed by connecting it to its internal knowledge to provide factual information. By isolating these steps and examining where models might fall short, we aim to pinpoint the layer-by-layer flow of information and advance our understanding about why models struggle to draw on their internal knowledge when relying on visual inputs.
Preliminaries and Notation.
VLMs are architectures that integrate visual and textual modalities to perform tasks such as image captioning, visual question answering and more. Our VLM analysis in this work is mainly focused on the common architecture of LLaVA Liu et al. (2023b), and thus, we will follow the same notation.
The VLM is composed of a pre-trained visual encoder $g$, a projection layer $W$, and an LLM $f$. Given an input image $\mathbf{X}_{\mathbf{v}}$ and text query $\mathbf{X}_{\mathbf{t}}$, $\mathbf{X}_{\mathbf{v}}$ is first processed by $g$ to provide the visual features $\mathbf{Z}_{\mathbf{v}}=g(\mathbf{X}_{\mathbf{v}})$ then passed through the projection layer to align the visual features into the LLM’s embedding space as a series of tokens $\mathbf{H}_{\mathbf{v}}=W\cdot\mathbf{Z}_{\mathbf{v}}$. In parallel, $\mathbf{X}_{\mathbf{t}}$ is passed through the LLM’s tokenizer and embedding layer to create a series of tokens $\mathbf{H}_{\mathbf{t}}$. Next the two are concatenated and passed through the LLM $f([\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{t}}])$. Figure 2, which illustrates the experiments we perform in the paper (see more details in Section 5 and 6), can be used also to better understand these notations.
Every token in this input can be written as $\mathbf{h}_{i,v}^{0}$ or $\mathbf{h}_{i,t}^{0}$ where the $v$ and $t$ indicate a visual or textual token, the $i$ indicates the $i^{\text{th}}$ visual or textual tokens, and the 00 indicates input to layer 00. As the language model processes the entire input at each layer $l$, it builds a representation $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell},\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}\right)$, where $\mathbf{h}_{1,g}^{\ell}...\mathbf{h}_{k,g}^{\ell}$ are the generated tokens hidden representations, $\left|\mathbf{H}_{\mathbf{v}}\right|=n$, $\left|\mathbf{H}_{\mathbf{t}}\right|=m$ and $\ell$ is the layer. We refer to the tokens once inside the LLM as hidden states.
In our experiments we focus on how information flows from the visual hidden states $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)$ to the textual and generated tokens. We do this by applying techniques of mechanistic interpretability.
Figure 2: The left pane is an illustration for the experiment in Section 5. The hidden states at the positions of the visual tokens are copied from the forward pass of the injected entity at the source layer (in this case layer 1) replacing the hidden states of the original entity’s forward pass at the same layer and position. The right pane is an illustration of the experiment in Section 6. Starting from the source layer, the hidden states at the positions of the visual tokens are copied and patched instead of the hidden states at the same positions in the subsequent layers until the target layer, effectively "freezing" their processing by the LLM.
Dataset and Experiments Overview.
To analyze how VLMs integrate visual and textual information, we introduce a new dataset called POPVQA in Section 3, featuring popular entities that VLMs are likely to have internal knowledge about. This dataset focuses on a two-hop reasoning process: (i) identification of an entity based on an image, followed by (ii) extraction of factual information related to that entity. We compare model responses when the entity is given as text (“Robin Williams”) versus as an image, providing insight into the gap between visual and textual representations. In these experiments, we first present a motivational case study in Section 4, demonstrating that various VLMs (e.g., LLaVA7B, Qwen2-VL) show a marked accuracy drop when relying on visual inputs. This experiment identifies a gap between the use of visual and language information in VLMs.
Motivated by the first experiment, we turn to further explore how VLMs rely on visual and textual information. We delve deeper into the reasons behind the gap, focusing on LLaVA with different encoders such as CLIP and SigLIP. Specifically, we aim to understand at what point in the inference pass the entity in an image is identified, and this identity is propagated to the query and generated token positions.
Through activation patching (Vig et al., 2020) and other mechanistic interpretability methods, we trace how and where entity recognition occurs, and how it impacts the model’s ability to retrieve the correct facts. Section 5 describes how we employ activation patching to identify the layer at which the model extracts the entity identity, establishing an upper bound on identification layers. Building on these findings, Section 6 applies forward activation patching to investigate whether the entity identity could be inferred at earlier layers.
We include additional experiments using attention knockouts (Geva et al., 2023) in appendix B, which further validate our conclusions. Throughout the paper, we use POPVQA to systematically evaluate VLM performance on visually driven factual retrieval and demonstrate how visual and textual components interact within a unified transformer architecture.
3 The POPVQA Dataset
Figure 3: The POPVQA dataset contains various entity-image pairs of celebrities, landmarks, brand logos and painting. For each entity we keep the image, its subject, popularity score, possible aliases and questions and answers related to that entity.
In order to isolate the model’s ability to extract knowledge about a given entity from its ability to identify it in the first place, we want to restrict our questions to entities that the model recognizes. Other datasets that have knowledge seeking goals such as infoSeek (Chen et al., 2023) or VeQuAE (Lerner et al., 2022) do not include the entity in the image but rather only the answer to the question, thus, making it difficult to do this separation. Therefore, we built the POPVQA dataset to facilitate this need. It is inspired by PopQA Mallen et al. (2023) that created a QA dataset for various entities using Wikipedia.
The first part of POPVQA comprises of 15,395 entity-image pairs from Wikidata (Vrandečić and Krötzsch, 2014) along with aliases and popularity score (number of page views). The vast majority of entities are celebrities (63.6%). To allow ablation of other entity types, we added landmarks (18%), such as famous structures or UNESCO sites, famous paintings (14.6%), and brand logos (3.8%). All images are resized to 336 × 336 which is a standard input resolution, by adding black borders to create a square, then resizing. This was done in order to maintain the original aspect ratio, without losing any details. The popularity score was used only in the creation of the dataset such that we select only entities with high popularity, which increases the chance that the models will recognize them Mallen et al. (2023). We do not use this value directly in the experiments.
@@ 3 The POPVQA Dataset
One legitimate concern is the presence of stale information, particularly for entities like celebrities whose personal circumstances (e.g., marital status, offspring) may change over time. However, this issue affects only a small subset of the entities in POPVQA, and an even smaller fraction of the associated questions. To mitigate this, we allow multiple correct answers per question where applicable. Additionally, we provide code for rebuilding POPVQA using up-to-date snapshots of Wikipedia and Wikidata, enabling users to generate a version that reflects more recent factual knowledge.
In the textual representation case, the model is presented with the question as is. For the visual case, the model receives an image of the entity along with a reformulated query, such as “Who is the spouse of the subject of this image?”. This ensures that the model needs to internally identify the subject depicted in the image before reasoning and generating an answer.
4 Motivating Experiment
We start by quantifying the model’s performance gap in entity knowledge extraction when presented with visual versus textual representations of entities. Given that the model recognizes the entity, asking about the visual representation, is not a harder task in essence, but a matter of computational capacity. Our motivating experiment demonstrate that, indeed, using the visual representation is more challenging. The subsequent experiments aim at explaining this gap and suggest that the reason for this bigger challenge is that the model needs to use part of its capacity to first identify the subject in the image and only then it can answer the question with the remaining capacity. Thus, as the model is of a fixed depth, once the model identifies the entity, it becomes a question of extracting the answer within the limits of the ‘remaining compute’ of the model.
Given the above discussion, we relate to the textual representation accuracy as the potential of the model. The lower the drop between the textual and visual accuracy, the better the model is at utilizing the visual input to elicit extraction of its internal knowledge. This could indicate better alignment between the embedding spaces of the LLM and the visual backbone.
To run the experiment, we first identify which entities each model is capable of recognizing. This is done by presenting the model with an image of each entity along with a prompt asking it to identify the subject (e.g., “Identify by name the subject in this image.”). Entities for which the model returns the correct name are considered identified. We then evaluate the model’s ability to answer factual questions for these identified entities given the two representations, textual versus visual, using the formulation in Section 3. This setup ensures that performance differences are not due to failures in entity recognition, but reflect differences in how well the model retrieves knowledge once the entity is known.
We show results for the following models; LLaVA-1.5-7B (Liu et al., 2023b) as the canonic base VLM; we refer to it as LLaVA7B. LLaVA-MORE-8B (Cocchi et al., 2025) which uses a larger Llama 3.1 as its LLM and we try both variants using CLIP (Radford et al., 2021b) and SigLIP (Zhai et al., 2023) as backbones allowing for controlled comparison between visual backbones; we refer to these as LM-CLIP and LM-SigLIP. We use LLaVA-1.6-34B to assess the effect of model scale, and refer to it as LLaVA34B. We also compare to Qwen2-VL (Wang et al., 2024) that has a different architecture and training scheme.
@@ Results.
Table 1 shows the results for the different models. We measured accuracy per entity as the proportion of correctly answered questions for that entity and calculated the overall mean accuracy across all entities. To verify the significance of the results, we performed a Wilcoxon signed-rank test between the per-entity accuracies for visual versus textual representations. The p-values range from $p=4.02\times 10^{-29}$ for the lower drop to $p=3.83\times 10^{-76}$ for the higher drop. A more fine grained analysis by the entity types can be found in appendix A.
Model #Identified Img Acc Txt Acc Drop
LLaVA7B 925 0.276 0.453 0.177
LM-CLIP 844 0.216 0.378 0.163
LM-SigLIP 660 0.201 0.377 0.176
LLaVA34B 1286 0.534 0.656 0.121
Qwen2-VL 3143 0.433 0.476 0.043
Table 1: Results on POPVQA. Reported accuracy is first calculated per subject then averaged across the entire dataset.
Surprisingly, LLaVA7B outperforms both LLaVA-MORE-8B variants, recognizing more entities and reaching higher accuracy. Both LLaVA-MORE-8B variants have similar accuracy, both with visual and textual representations, but the variant that uses SigLIP as a visual backbone identifies considerably more entities. LLaVA34B has significantly better results with a lower drop showing the benefit of increasing model depth to improve the computational capacity. Qwen2-VL also exhibits strong results, with the lowest drop, pointing to the advantage of more modern training schemes which include unfreezing the visual backbone parameters during training.
@@
Interestingly, a common mistake made by the model when asked to name the subject’s spouse/parent/sibling/child, is to name the subject itself. This observation provides an indication to our hypothesis that by the time information about the entity propagates from the image tokens to later token positions, there is not enough computational capacity left to use this information, and therefore the model outputs the entity’s name.
@@
Note, though, that some entities did exhibit an increase in accuracy with visual inputs. This occurred specifically when textual input alone was insufficient for reasoning, but visual input provided contextually inferable details. For example, a subject wearing the French national team uniform, implicitly informed the model that the language spoken by the subject is French.
5 Late Information Flow
To analyze the information flow in the VLMs, we apply interventions in their inference. We continue to focus on entity-image pairs for which, without intervention, the model correctly identifies the entity in the image. For all the following experiments we focus on the LLaVA architecture due to its simple architecture that allows for controlled probing, leaving other architectures as future work. We use LLaVA7B which successfully recognizes 925 entities from POPVQA. Yet, in order to show comparative results with different backbones we show also results on LM-SigLIP and LM-CLIP as well. These backbones were chosen because their training process inherently brings visual and textual representations closer which makes them a common choice in VLMs.
@@ 5 Late Information Flow
We begin by identifying the point of critical information flow from the image hidden representation to later positions. This will provide an upper bound on the layer depth at which entity identification happens and point us to where this information is likely to propagate onward. We can subsequently infer that by this point the model mostly finalizes its image-based representation enrichment and focuses on leveraging this information for prediction. We hypothesize that this happens in deep layers, leaving few layers for factual knowledge extraction, which may begin to explain the performance gap presented in Section 4.
@@
The experiment involves two distinct forward passes through the model, using different entities: an original entity and a injected entity (as shown in Figure 2). Each forward pass is initiated with a prompt asking the model to identify the subject of an image. We extract the image-token representations $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$ from the injected entity forward pass at a specific layer. These representations are then patched at the same token positions and same layer of the forward pass of the original entity. Hence, we call this technique cross patching. After patching, we allow the model to complete its forward pass and predict which entity is identified. The layer at which the predicted entity switches from being the injected entity to the original entity identifies the critical point where the model has completed sufficient processing of the visual representation such that patching does not affect its prediction. When the patching happens between entities of the same type, we use the original prompt. When we patch between different type entities, we replace the reference word (painting, place, etc.) with the word "subject" so as not to bias the output to a certain type.